# Low-Rank Bandits via Tight Two-to-Infinity Singular Subspace Recovery

### v1

@@ 6.3 Sample complexity lower bounds
Let $\delta\in(0,1/2)$ and assume that $\xi_{t}\sim\mathcal{N}\left(0,1\right)$. The sample complexity of any $(\epsilon,\delta)$-PAC estimator of $v^{\pi}$ must satisfy
$T\geq\frac{L_{M,\pi}}{2\varepsilon^{2}}\operatorname{kl}\left(\delta,1-\delta\right)$
@@
Let $\delta\in(0,1/2)$ and assume that $\xi_{t}\sim\mathcal{N}\left(0,1\right)$. There exists a target policy $\pi$ and a homogeneous matrix $M\in{\cal M}(c,c^{\prime})$ such that the sample complexity of any $\left(\varepsilon,\delta\right)-$PAC estimator of $v^{\pi}$ must satisfy
$T\gtrsim\frac{m+n}{\omega_{\max}mn\varepsilon^{2}}\operatorname{kl}\left(\delta,1-\delta\right).$
Note that when $\omega$ is homogeneous, the sample complexity lower bound reduces to $T\gtrsim\frac{\left(m+n\right)}{\varepsilon^{2}}\operatorname{kl}\left(\delta,1-\delta\right)$. Since $\operatorname{kl}\left(\delta,1-\delta\right)\geq\log\left(1/\left(2.4\delta\right)\right)$ (Kaufmann et al., 2016), this lower bound nearly matches the sample complexity guarantees of SIPS and RS-PE when $M$ and $\omega$ are homogeneous.
@@ 7 Best Policy Identification
As for the PE learning task, we present two BPI algorithms, the first directly based on the estimate $\widehat{M}$, and the second following our two-phase approach. These algorithms, referred to as SBPI and RS-BPI, output the policy $\hat{\pi}_{\text{\tt SBPI}}$ and $\hat{\pi}_{\text{\tt RS-BPI}}$ respectively, defined by, for all context $i$,
@@
  Input: Budgets of rounds $T$ and $T_{1}$, context distribution $\rho$, regularization parameter $\tau$, confidence level $\delta$,
@@
Let $\varepsilon,\delta\in(0,1)$ and assume that $\rho$ is homogeneous.
(i) SBPI is $(\varepsilon,\delta)$-PAC as soon as
$T\gtrsim\frac{\mu^{11}\kappa^{4}r^{3}(m+n)^{3}}{(m\wedge n)^{2}\varepsilon^{2}}\log^{3}\!\left(\frac{(m+n)T}{\delta}\right)\!.$
(ii) Choosing $\tau\leq r(m\wedge n)^{-1}(\rho_{\min}/\rho_{\max})\log(16d/\delta)$, and $T_{1}=\lfloor T/2\rfloor$, RS-BPI is $(\varepsilon,\delta)$-PAC as soon as
$T\gtrsim\!\frac{\mu^{2}r\left(m+n\right)}{\varepsilon^{2}}\log\left(\frac{mn}{\delta}\right)\!+O\!\left(\!{1\over\varepsilon}\!\right)\!\log^{3}\!\left(\frac{(m+n)T}{\delta}\right)\!,$
where the term $O({1\over\varepsilon})$ hides a multiplicative factor depending [^8] Refer to (37) in Appendix F for a complete expression. polynomially on $\mu,\kappa,r,m,n$.
Note that under our homogeneity assumption on $\rho$, the sample complexity of RS-BPI  scales as $\frac{\mu^{2}r\left(m+n\right)}{\varepsilon^{2}}\log\left(1/{\delta}\right)$ when $\varepsilon$ is sufficiently small. In this case, it exhibits a better dependence on the model parameters than the sample complexity of SBPI when $M$ is not homogeneous. For large $\varepsilon$, the guarantees of SBPI can be better than those of RS-BPI. Nonetheless, our experiments (see Appendix J.2) suggest that RS-BPI still outperforms SBPI in the large $\varepsilon$ regime.
A known minimax sample complexity lower bound for BPI in contextual low-rank bandits scales as ${r(m+n)\over\varepsilon^{2}}\log(1/\delta)$ (this is essentially proved by (Lee et al., 2023) for lumpable bandits, a subclass of low-rank bandits). Hence, in view of Theorem 7.1, both SBPI and RS-BPI are nearly minimax optimal when the reward matrix and context distribution are homogeneous; the sample complexity of SBPI has an additional $\log^{2}(1/\delta)$ term, and RS-BPI is minimax optimal when $\varepsilon$ is sufficiently small.
Finally, it is worth noting that for the two proposed algorithms, data is just gathered using the uniform sampling policy (no need to be adaptive to achieve a minimax optimal sample complexity when $M$ and $\rho$ are homogeneous).
@@ 8 Regret Minimization
In this section, we introduce RS-RMIN, our algorithm designed for regret minimization in contextual low-rank bandits. Presented as Algorithm 2, RS-RMIN runs in two phases. In the first phase, we collect $T_{1}$ samples by just selecting arms uniformly at random. Based on the corresponding reward observations, we estimate the singular subspaces of $M$ using the procedure outlined in Subsection 4.1. The problem is then reformulated as a misspecified contextual linear bandit as shown in Section 5.
@@
$\beta(\delta)=\sqrt{2\log\left(\frac{T_{2}mn}{\delta}\left\lceil\frac{1}{2}\log\left(\frac{T_{2}}{d}\right)\right\rceil\right)}+1.$ (16)
@@
$f(T,M,\rho):=\mu^{3}\kappa^{2}r^{7/4}\frac{(m+n)^{3/4}(mn)^{1/4}}{\sqrt{m\min_{i}\rho_{i}}(m\wedge n)^{1/2}}\sqrt{T}.$
@@
Here, $\widetilde{\Theta}(\cdot),\widetilde{O}(\cdot)$ may hide logarithmic factors in $T$, $m$ and $n$. Theorem 8.1 is proved in Appendix G. When the reward matrix $M$ and context distribution $\rho$ are homogeneous, the regret upper bound of RS-RMIN simplifies to:
$R^{\pi}(T)=\widetilde{O}\left((m+n)^{3/4}\sqrt{T}\right).$
Observe that if we were to use an algorithm meant for contextual bandits with no structure, then the best regret guarantee we can obtain is $\widetilde{O}(\sqrt{mnT})$ (Lattimore & Szepesvári, 2020). In contrast, our regret upper bound scales as $\widetilde{O}((m+n)^{3/4}\sqrt{T})$, and clearly $(m+n)^{3/4}\sqrt{T}\ll(m+n)\sqrt{T}=\Theta(\sqrt{mnT})$ for all $T\geq 1$ whenever $m=\Theta(n)$. In other words, our algorithm truly takes advantage of the low-rank structure. We finally remark that Lee et al. (2023) also provides non-trivial regret bounds for general low rank reward matrices, as discussed in the introduction, but only when $T=\widetilde{O}((m+n)^{2(1+(3r)^{-1})})$.
@@
  Input: Budget of rounds $T$, $T_{1}$, context distribution $\rho$
@@ 9 Conclusion
For PE and BPI, we follow such a two-stage approach and propose RS-PE and RS-BPI whose sample complexity enjoy, in the worst case, a scaling of order $\widetilde{O}(r(m+n)/\varepsilon^{2})$. We also considered a single-stage approach to propose SIPS and SBPI with a similar worst-case sample complexity. However, experimentally, RS-PE and RS-BPI exhibit far superior performance in comparison with SIPS and SBPI. Shedding some light on this discrepancy is of future interest.
For regret minimization, the benefits of the two-stage approach are even more pronounced. Indeed, the proposed RS-RMIN  achieves a non-trivial regret upper bound scaling as $\widetilde{O}((m+n)^{3/4}\sqrt{T})$. It is unclear if achieving $\sqrt{(m+n)T}$ is possible for general low-rank reward matrices, but we believe this to be an exciting research question.
Impact statements
@@ Appendix A Additional Related Work
In this Appendix, we provide further additional work that is related to our work. Most specifically, we survey some of the recent algorithms for misspecified contextual linear bandits. Next, we also discuss bilinear bandits from which the idea of reducing a low-rank bandit problem to an almost low-dimensional linear bandit originated.
@@
Misspecified contextual linear bandits have recently received a lot of attention (Gopalan et al., 2016; Ghosh et al., 2017; Lattimore et al., 2020; Foster et al., 2020; Zanette et al., 2020; Takemura et al., 2021). The best achievable minimax regret bounds for this setting are $\widetilde{O}(d\sqrt{T}+\epsilon\sqrt{d}T)$ where $d$ is the ambient dimension, $\epsilon$ is the misspecification level, and $T$ is the time horizon. This guarantee is achievable even if $\epsilon$ is not known as proven by (Foster et al., 2020). Interestingly, (Lattimore et al., 2020) show that one cannot hope to improve the dependence on $d$ in the term $\epsilon\sqrt{d}$ due to misspecification for most interesting regimes. However, the first term, that is of order $d\sqrt{T}$, is due to the assumption that the set of arms per context is continuous or infinite. In fact, (Lattimore et al., 2020) also shows that a minimax regret of order $\widetilde{O}(\sqrt{dT\log(K)}+\epsilon\sqrt{d}T)$ in the case of linear bandit when the number of arms is $K$. Indeed, in the case of misspecified contextual linear bandits when the number of arms per context is finite, say $K$, (Takemura et al., 2021) show that SupLin-UCB, initially proposed by (Chu et al., 2011), achieves a regret guarantee of order $\widetilde{O}(\sqrt{d\log\phantom{)\!\!\!}^{2}(K)T}+\epsilon\sqrt{d\log(K)}T)$ without knowledge of $\epsilon$. They further propose a variant of SupLin-UCB that has an improved regret upper bound of order $\widetilde{O}(\sqrt{d\log(K)T}+\epsilon\sqrt{d}T)$.
@@
In this setting, originally introduced by (Jun et al., 2019), in each round, the learner selects a pair of features $(x,y)$ in $\mathcal{X}\times\mathcal{Y}\subseteq\mathbb{R}^{m}\times\mathbb{R}^{n}$, and then observes a reward with expected value $x^{\top}My$, where $M$ is assumed to be of rank $r\ll m+n$. This setting is closely related to ours, because if one assumes that $\mathcal{X}$ and $\mathcal{Y}$ correspond to the canonical basis in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$, respectively, then one recovers our reward model. To solve this problem, (Jun et al., 2019) proposed ESTR, a two-stage algorithm, where the algorithm first estimates the singular subspaces of $M$ using spectral methods, and then solves an almost low-dimensional linear bandit problem, using an adaptation of OFUL (Abbasi-Yadkori et al., 2011) referred to as LowOFUL. The key insight is that subspace recovery enables a reformulation of bilinear bandit problems as low-dimensional linear bandits. This lead to the first regret guarantees of order $\widetilde{O}((m+n)^{3/2}\sqrt{T})$. Since then, there has been many subsequent work with new algorithms, improvements, and various generalizations (Jang et al., 2021; Lu et al., 2021; Kang et al., 2022; Cai et al., 2023). Notably, (Kang et al., 2022) proposes an algorithm, G-ESTT, to solve generalized bilinear bandit problems. Their algorithm achieves the state-of-the-art regret guarantee of order $\widetilde{O}((m+n)\sqrt{T})$, and relies on the same two-stage algorithm design idea as (Jun et al., 2019), but uses instead a novel subspace estimator that is based Stein’s method. A major limitation of existing state-of-the-art algorithms for bilinear bandits is that when $|\mathcal{X}\times\mathcal{Y}|=\widetilde{O}(mn)$, as this is the the case in our settings, then the obtained regret guarantees are no better than the ones achievable by minimax optimal algorithms for unstructured bandits. Indeed, an algorithm like UCB that does not use the low-rank structure achieves a regret rate of order $\widetilde{O}(\sqrt{mnT})$ (Lattimore & Szepesvári, 2020)).

### v2

@@ 6.3 Sample Complexity Lower Bounds
Let $\varepsilon>0,\delta\in(0,1/2)$ and assume that $\xi_{t}\sim\mathcal{N}\left(0,\sigma^{2}\right)$. The sample complexity of any $(\epsilon,\delta)$-PAC estimator of $v^{\pi}$ must satisfy
$T\geq\frac{\sigma^{2}L_{M,\pi}}{2\varepsilon^{2}}\operatorname{kl}\left(\delta,1-\delta\right)$
@@
Let $\varepsilon>0,\delta\in(0,1/2)$ and assume that $\xi_{t}\sim\mathcal{N}\left(0,\sigma^{2}\right)$. There exists a target policy $\pi$ and a homogeneous matrix $M\in{\cal M}(c,c^{\prime})$ such that the sample complexity of any $\left(\varepsilon,\delta\right)-$PAC estimator of $v^{\pi}$ must satisfy
$T\gtrsim\frac{\sigma^{2}\left(m+n\right)}{\omega_{\max}mn\varepsilon^{2}}\operatorname{kl}\left(\delta,1-\delta\right).$
Note that when $\omega$ is homogeneous, the sample complexity lower bound reduces to $T\gtrsim\frac{\sigma^{2}\left(m+n\right)}{\varepsilon^{2}}\operatorname{kl}\left(\delta,1-\delta\right)$. Since $\operatorname{kl}\left(\delta,1-\delta\right)\geq\log\left(1/\left(2.4\delta\right)\right)$ (Kaufmann et al., 2016), this lower bound nearly matches the sample complexity guarantees of RS-PE when $M$ and $\omega$ are homogeneous. If in addition, $\sigma\gtrsim\|M\|_{\max},$ it also matches the sample complexity guarantees of SIPS.
@@ 7 Best Policy Identification
As for the PE learning task, our first algorithm is directly based on the estimate $\widehat{M}$, and the second follows our two-phase approach. These algorithms, referred to as SBPI and RS-BPI, output the policy $\hat{\pi}_{\text{\tt SBPI}}$ and $\hat{\pi}_{\text{\tt RS-BPI}}$ respectively, defined by, for all $i\in[m]$,
@@
  Input: Budgets of rounds $T$ and $T_{1}$, context distribution $\rho$, regularization parameter $\tau$
@@
Let $\varepsilon\in\left(0,\|M\|_{\max}\right)$, $\delta\in\left(0,1\right)$ and assume that $\rho$ is homogeneous.
(i) $v^{\star}-v^{\hat{\pi}_{\text{\tt SBPI}}}\leq\varepsilon$ with probability larger than $1-\delta$ as soon as
$T\gtrsim\frac{L^{2}\mu^{6}\kappa^{4}r^{3}(m+n)^{3}}{(m\wedge n)^{2}\varepsilon^{2}}\log^{3}\!\left(\frac{(m+n)T}{\delta}\right)\!.$
(ii) Choosing $\tau\leq r(m\wedge n)^{-1}(\rho_{\min}/\rho_{\max})\log(16d/\delta)$ and $T_{1}=\lfloor T/2\rfloor$, $v^{\star}-v^{\hat{\pi}_{\text{\tt RS-BPI}}}\leq\varepsilon$ with probability larger than $1-\delta$ as soon as
$T\gtrsim\frac{\sigma^{2}\mu^{2}r\!\left(m+n\right)\!}{\varepsilon^{2}}\log\!\left(\!\frac{m+n}{\delta}\!\right)\!+\!\frac{K}{\varepsilon}\log^{3}\!\left(\!\frac{(m+n)T}{\delta}\!\right)\!,$
where $K$ depends polynomially on the model parameters  [^8] Refer to (44) in Appendix F for a complete expression. $\|M\|_{\max},\sigma,\mu,\kappa,r,m,n$.
Note that under our homogeneity assumption on $\rho$, the sample complexity guarantees of RS-BPI  scale as $\frac{\sigma^{2}\mu^{2}r\left(m+n\right)}{\varepsilon^{2}}\log\left(1/{\delta}\right)$ when $\varepsilon$ is sufficiently small. In this case, it exhibits a better dependence on the model parameters than the sample complexity guarantees of SBPI when $M$ is not homogeneous. For large $\varepsilon$, the guarantees of SBPI can be better than those of RS-BPI. Nonetheless, our experiments (see Appendix J.2) suggest that RS-BPI still outperforms SBPI in the large $\varepsilon$ regime.
A known minimax sample complexity lower bound for BPI in contextual low-rank bandits scales as ${r(m+n)\over\varepsilon^{2}}\log(1/\delta)$ (this is essentially proved by (Lee et al., 2023) for context-lumpable bandits, a subclass of contextual low-rank bandits). Hence, in view of Theorem 7.1, both SBPI and RS-BPI are nearly minimax optimal when the reward matrix and context distribution are homogeneous; the sample complexity of SBPI has an additional $\log^{2}(1/\delta)$ term, and RS-BPI is minimax optimal when $\varepsilon$ is sufficiently small.
Finally, it is worth noting that for the two proposed algorithms, data is simply gathered using the uniform sampling policy (there is no need to be adaptive to achieve a minimax optimal sample complexity when $M$ and $\rho$ are homogeneous).
@@ 8 Regret Minimization
In this section, we introduce RS-RMIN, our algorithm designed for regret minimization in contextual low-rank bandits. Presented as Algorithm 2, RS-RMIN runs in two phases. In the first phase, we collect $T_{1}$ samples by simply selecting arms uniformly at random. Based on the corresponding reward observations, we estimate the singular subspaces of $M$ using the procedure outlined in Subsection 4.1. The problem is then reformulated as a misspecified contextual linear bandit, as described in Section 5.
@@
$\beta(\delta)=\sigma(1+\sqrt{2\log\left(\frac{T_{2}mn}{\delta}\left\lceil\frac{1}{2}\log\left(\frac{T_{2}}{d}\right)\right\rceil\right)}).$ (12)
@@
$f(T,M,\rho):=L\mu^{2}\kappa^{2}r^{5/4}\frac{(m+n)^{3/4}(mn)^{1/4}}{\sqrt{m\rho_{\min}}(m\wedge n)^{1/2}}\sqrt{T}.$
@@
Here, $\widetilde{\Theta}(\cdot),\widetilde{O}(\cdot)$ may hide logarithmic factors in $T$, $m$, $n$ and $L/\sigma$. Theorem 8.1 is proved in Appendix G. When the reward matrix $M$ and context distribution $\rho$ are homogeneous, the regret upper bound of RS-RMIN simplifies to:
$R^{\pi}(T)=\widetilde{O}\left(Lr^{5/4}n^{3/4}\sqrt{T}\right).$
Observe that if we were to use an algorithm meant for contextual bandits with no structure, then the best regret guarantee we can obtain is $\widetilde{O}(\sqrt{mnT})$ (Lattimore & Szepesvári, 2020). In contrast, our regret upper bound scales as $\widetilde{O}(n^{3/4}\sqrt{T})$, and clearly $n^{3/4}\sqrt{T}\ll n\sqrt{T}=\Theta(\sqrt{mnT})$ for all $T\geq 1$ whenever $m=\Theta(n)$. In other words, our algorithm truly takes advantage of the low-rank structure. We finally remark that Lee et al. (2023) also provides non-trivial regret bounds for general low-rank reward matrices, as discussed in the introduction, but only when $T=\widetilde{O}((m+n)^{2(1+(3r)^{-1})})$.
@@
  Input: Budget of rounds $T$ and $T_{1}$, context distribution $\rho$
@@
We also explain in Appendix H why reducing the problem to almost-low-dimensional linear bandits, a framework employed for regret minimization in previous works such as (Jun et al., 2019; Lu et al., 2021; Kang et al., 2022), gives suboptimal regret guarantees compared to our reduction to misspecified linear bandits. Additionally, we present an algorithm based on the reduction to almost-low-dimensional linear bandits that leverages our entrywise guarantees, with regret $R^{\pi}(T)=\widetilde{O}((m+n)\sqrt{T})$ in the homogeneous case. Our results complement observations made in (Kang et al., 2024), where reducing to misspecified linear bandits yields $\Theta(r^{1/4}(m+n)^{1/4})$ higher regret than reducing to almost-low-dimensional linear bandits.
As mentioned in Section 4.2, using spectral methods introduces in the subspace recovery error bound a dependence on $L=\|M\|_{\max}\vee\sigma$ instead of $\sigma$ (see also (Keshavan et al., 2009)). This may affect the regret bounds one may obtain in settings where $\sigma\ll\|M\|_{\max}$. Indeed, should this improvement be possible, one may expect a regret upper bound of order $\widetilde{O}(\sigma(n+m)^{3/4}\sqrt{T}+\|M\|_{\max}(n+m))$ when $M$ and $\omega$ are homogeneous (see Appendix I.4 for more details) which is better than ours for $\sigma\ll\|M\|_{\max}$. We believe that the dependence on $\sigma$ can be improved using either the method proposed in Appendix I.3 or methods based on nuclear norm regularization (Chen et al., 2020) as discussed in Appendix I.4.
Proof sketch of Theorem 8.1.
For the sake of simplicity, we assume that $\mu,\kappa=\Theta(1)$ and $m=\Theta(n)$ in the proof sketch. The regret accumulated during the exploration phase (Phase 1) is upper bounded by $2T_{1}\|M\|_{\max}$. To upper bound the regret of the second phase, we observe that we face a misspecified linear bandit with misspecification $\epsilon_{\max}=\widetilde{O}\left((L^{2}r)/(T_{1}\sigma_{r}\omega_{\min})\right)$ for $T_{1}$ sufficiently large according to Corollary 5.1, and we apply the regret bound for misspecified linear bandits from Theorem 1 in (Takemura et al., 2021) (see Theorem G.1). Putting all of this together, we get:
$R^{\pi}(T)=\widetilde{O}\left(T_{1}\|M\|_{\max}+\frac{L^{2}}{T_{1}}\frac{r^{3/2}\sqrt{n}T}{\sigma_{r}\omega_{\min}}+L\sqrt{rnT}\right).$
Choosing $T_{1}$ that minimizes this expression while using that $\|M\|_{\max}/\sigma_{r}\lesssim r/\sqrt{mn}$ (see Lemma B.8) yields the claimed regret.
@@ 9 Conclusion
For PE and BPI, we followed such a two-stage approach and proposed RS-PE and RS-BPI whose minimax sample complexity typically scales as $\widetilde{O}(r(m+n)/\varepsilon^{2})$. We also considered a single-stage approach to propose SIPS and SBPI with a similar minimax sample complexity. However, experimentally, RS-PE and RS-BPI exhibit far superior performance in comparison with SIPS and SBPI. Shedding some light on this discrepancy is of future interest.
For regret minimization, the benefits of the two-stage approach are even more pronounced. Indeed, the proposed RS-RMIN  achieves a non-trivial regret upper bound typically scaling as $\widetilde{O}(r^{5/4}(m+n)^{3/4}\sqrt{T})$. It is unclear if achieving $\widetilde{O}(\sqrt{\mathrm{poly}(r)(m+n)T})$ is possible for general low-rank reward matrices, but we believe this to be an exciting research question.
Acknowledgements
This research was supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, the Swedish Research Council (VR), and Digital Futures.
Impact Statement
@@ Appendix A Additional Related Work
In this appendix, we provide additional related work. Most specifically, we survey some of the recent algorithms for misspecified contextual linear bandits. Next, we also discuss bilinear bandits from which the idea of reducing a low-rank bandit problem to an almost low-dimensional linear bandit originated.
@@
Misspecified contextual linear bandits have recently received a lot of attention (Gopalan et al., 2016; Ghosh et al., 2017; Lattimore et al., 2020; Foster et al., 2020; Zanette et al., 2020; Takemura et al., 2021). The best achievable minimax regret bounds for this setting are $\widetilde{O}(d\sqrt{T}+\epsilon\sqrt{d}T)$ where $d$ is the ambient dimension, $\epsilon$ is the misspecification, and $T$ is the time horizon. This guarantee is achievable even if $\epsilon$ is not known as proven by (Foster et al., 2020). Interestingly, (Lattimore et al., 2020) show that one cannot hope to improve the dependence on $d$ in the term $\epsilon\sqrt{d}$ due to misspecification for most interesting regimes. However, the first term, that is of order $d\sqrt{T}$, is due to the assumption that the set of arms per context is continuous or infinite. In fact, (Lattimore et al., 2020) also shows that a minimax regret of order $\widetilde{O}(\sqrt{d\log(K)T}+\epsilon\sqrt{d}T)$ is achievable when the misspecified linear bandit has a finite number of arms $K$. Furthermore, in the case of misspecified contextual linear bandits with a finite number of arms per context, say $K$, (Takemura et al., 2021) show that SupLinUCB, initially proposed by (Chu et al., 2011), achieves a regret guarantee of order $\widetilde{O}(\sqrt{d\log\phantom{)\!\!\!}^{2}(K)T}+\epsilon\sqrt{d\log(K)}T)$ without knowledge of $\epsilon$. They further propose a variant of SupLinUCB that has an improved regret upper bound of order $\widetilde{O}(\sqrt{d\log(K)T}+\epsilon\sqrt{d}T)$.
@@
In this setting, originally introduced by (Jun et al., 2019), in each round, the learner selects a pair of features $(x,y)$ in $\mathcal{X}\times\mathcal{Y}\subseteq\mathbb{R}^{m}\times\mathbb{R}^{n}$, and then observes a reward with expected value $x^{\top}My$, where $M$ is assumed to be of rank $r\ll m+n$. This setting is closely related to ours, because if one assumes that $\mathcal{X}$ and $\mathcal{Y}$ correspond to the canonical basis in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$, respectively, then one recovers our reward model. To solve this problem, (Jun et al., 2019) proposed ESTR, a two-stage algorithm that first estimates the singular subspaces of $M$ using spectral methods, and then solves an almost low-dimensional linear bandit problem, using an adaptation of OFUL (Abbasi-Yadkori et al., 2011) referred to as LowOFUL. The key insight is that subspace recovery enables a reformulation of bilinear bandit problems as low-dimensional linear bandits. This lead to the first regret guarantees of order $\widetilde{O}((m+n)^{3/2}\sqrt{T})$. Since then, there has been many subsequent work with new algorithms, improvements, and various generalizations (Jang et al., 2021; Lu et al., 2021; Kang et al., 2022; Cai et al., 2023). Notably, (Kang et al., 2022) proposes an algorithm, G-ESTT, to solve generalized bilinear bandit problems. Their algorithm also achieves a regret guarantee of order $\widetilde{O}((m+n)^{3/2}\sqrt{T})$, and relies on the same two-stage algorithm design idea as (Jun et al., 2019), but uses instead a novel subspace estimator that is based on Stein’s method. A major limitation of existing state-of-the-art algorithms for bilinear bandits is that when $|\mathcal{X}\times\mathcal{Y}|=\widetilde{O}(mn)$, as is the case in our setting, the obtained regret guarantees are no better than the ones achievable by minimax optimal algorithms for unstructured bandits. Indeed, an algorithm such as UCB that does not use the low-rank structure achieves a regret rate of order $\widetilde{O}(\sqrt{mnT})$ (Lattimore & Szepesvári, 2020)).
It is worth noting that in a concurrent work, (Jang et al., 2024) derive a regret bound that scales with $\tilde{O}((m+n)^{1/4}\sqrt{BT})$ where $B$ is a constant that depends on the action set geometry. This constant appears in their bounds because they use an experimental design approach. It is unclear how this constant would scale with our action set, but for the case of the unit sphere, which would also include our action set, they obtain a regret bound of order $\widetilde{O}((m+n)^{5/4}\sqrt{T})$. This still does not attain a sub-linear dependence on $m+n$ as we do in our work.